# Maintaining Distributed Data Structures in Dynamic Peer-to-Peer Networks

### v1

@@ 2.5 Merge
1.
The leader $x_{0}$ checks node $v$’s ID/key in $\mathcal{C}$ on which $X$ is currently located;
2.
$x_{0}$ communicates $v$’s ID/key to its followers in $\mathcal{B}$, i.e., to the nodes in $X\setminus\{x_{0}\}$;
3.
$x_{0}$ computes the split $X^{\prime}=\{u\in V:u>z\}$ where $z$ is the right neighbor of $v$ at level $\ell$ and elects the element with smallest ID/key in $X^{\prime}$ as a leader of such new cohesive group;
4.
Each node $u\in\hat{X}$, where $\hat{X}=X\setminus X^{\prime}$, sends a message $m(u)=\langle v,z,\text{``down''},\ell\rangle$ to its children $\Gamma(u)$ in $\mathcal{B}$;
5.
Each node $w\in X^{\prime}$, sends a message $m(w)=\langle v,z,\text{``right''},\ell\rangle$ to its children $\Gamma(w)$ in $\mathcal{B}$;
6.
The new cohesive group $X^{\prime}$ separates from $X$ and continues its merge procedure on $z$ at level $\ell$;
7.
If the current level $\ell$ is at most $\ell_{\max}(x_{0})$, $\hat{X}$ must be merged between $v$ and $z$ in $\mathcal{C}$:
7.a.
$\hat{X}$ merges itself between $v$ and $z$;
7.b.
Each node $u\in\hat{X}$ sends a message $m(u)=\langle u,y,\text{``down''},\ell\rangle$ where $y$ is $u$’s right neighbor in the newly merged area that includes $v,\hat{X}$, and $z$ to $\Gamma(u)$;
7.c.
Each node in $u\in\hat{X}$ announces its successful merge at level $\ell$ to its children $\Gamma(u)$;
8.
If the next level $\ell-1\geq 0$, then $\hat{X}$ must proceed to the next level:
8.a.
If at least one of $\hat{X}$ neighbors at level $\ell$ is a node in $\mathcal{B}$, i.e., $\{v,z\}\cap\mathcal{B}\neq\emptyset$, then $\hat{X}$ waits for $\{v,z\}\cap\mathcal{B}$ to be merged at level $\ell-1$ and proceeds with its own merge at level $\ell-1$;
Furthermore, once a node is fully merged in level 00 of $\mathcal{C}$, it enters the done state and terminates. Next, we describe the behavior of the nodes in $\mathcal{B}$ upon receiving messages from their parents. A node $v$, in the buffer network $\mathcal{B}$ in idle state is continuously listening to its parents. $v$’s behavior changes according to the type of message it is receiving. More precisely, let us assume that $v$ receives a message of the type $m(u)=\langle L,R,\texttt{move},\texttt{level}\rangle$ from its parent $u$:
If the message is sent by a node that is in the merge state at some level above $\ell_{\max}(v)$, then $v$ has to update its pointers:
1.
If $v<m(u).R$ and $v.\texttt{starting\_point}\leq m(u).L$ then $v$ updates its starting point and level for its own merge with the node $m(u).L$ and $m(u).\texttt{level}$, respectively;
2.
If $v>m(u).R$ and $v.\texttt{starting\_point}<m(u).R$ then $v$ updates its starting point and level for its own merge with the node $m(u).R$ and $m(u).\texttt{level}$, respectively;
3.
@@
4.
[…]
4.a.
$v$ looks at its left neighbor $z$ at level $\ell_{\max}(v)$ in $\mathcal{B}$. If $z$’s starting point is the same as $v$’s one, then $v$ sets itself as follower, otherwise $v$ becomes leader;
4.b.
[…]
1.
$v$ looks at its left neighbor $z$ at level $\ell_{\max}(v)$ in $\mathcal{B}$. If $z$’s starting point is the same as $v$’s one, then $v$ sets itself as follower, otherwise $v$ becomes leader;
2.
[…]
First and foremost, we notice that all the idle nodes in $\mathcal{B}$ correctly update their starting points. More precisely, a node $v\in\mathcal{B}$ that updates its starting point, it can choose only another point laying in its optimal search path from the top-left sentry to $\mathcal{C}$’s bottom most level. This property follows by the fact that each node $v$ in $\mathcal{B}$ receives messages only by its (left and right) parents in $\mathcal{B}$, i.e., $v$ receives a list of nodes ID that lay on the optimal search paths of its parents in the skip list $\mathcal{C}^{\prime}=\mathcal{C}\cup\mathcal{B}_{>\ell}$ obtained by merging $\mathcal{C}$ and the elements of $\mathcal{B}$ situated in the level above $\ell$. Thus, $v$ updates its starting point with nodes that it would encounter had it performed a classic insertion in $\mathcal{C}^{\prime}$. Thus, $v$’s new starting point lays in its optimal search path. This property, allows the nodes in $\mathcal{B}$ to “virtually” travel in $\mathcal{C}$ using the information shared by their parents. Next, we notice that one critical part of our protocol is that a cohesive group $X$ that was successfully merged at some level $\ell$ and wants to proceed one level down to $\ell-1$ may have to wait for its parents $v$ and $z$ to be successfully merged in such target level. This happens if, at least one of $X$’s neighbors at level $\ell$ in $\mathcal{C}$ is an element in the buffer (i.e., it is one of its parents in $\mathcal{B}$) and it is still in its merging phase at level $\ell-1$. The next lemma provides a bound on the time a cohesive group $X$ needs to merge itself in the skip list provided that it may have to wait for its parents to be successfully merged beforehand at each level.
@@
Assume that the cohesive group $X$ was merged at level $\ell$ of the clean network and has to proceed with its own merge at level $\ell-1$. Let $v$ and $z$ be $X$’s left and right parents in $\mathcal{B}$. We have three cases to consider: (i) $X$ is independent from its parents; (ii) $X$ depends on at least one of its parents through all the levels $\ell^{\prime}<\ell$; and, (iii) $X$ may depend on one of its parents for some time and become independent at some point. In the first case, $X$ is merging itself in a sub-skip list that contains only elements in the clean network. Thus, $X$’s overall merging time requires $\mathcal{O}(\log n)$ w.h.p. To analyze the second case, we notice that as $X$ depends on its parents, also $X$’s parents may depend on their parents as well and so on. We can model the time a cohesive group $X$ at level $\ell$ has to wait in order to start its own merge procedure at level $\ell-1$ as the time it would have taken had it traveled in a pipelined fashion with its parents. Moreover, such pipelining effect can be extended to include, in a recursive fashion, the parents of $X$’s parents, the parents of the parents of $X$’s parents, and so on. Thus, $X$ can be modeled as the tail of such “parents-of-parents” chain (or pipeline), in which $X$ must wait for all the nodes in the pipeline to be merged at level $\ell-1$ before proceeding with its own merge in such level. Moreover, we observe that, by our assumption, $X$ is dependent on at least one of its parents throughout the merge phase. Without loss of generality, let us assume that $X$ depends on one parent, say $z$. In other words, $X$ is never splitting, and its placing itself right after/before $z$ at each level $\ell^{\prime}<\ell$. The time $X$ takes to merge in one level $\ell^{\prime}$ after $z$ is $\mathcal{O}(1)$ (see Lemma 2.12). Thus, $X$ merge itself in some level $\ell^{\prime}$ by time $T_{z}^{\ell^{\prime}}+\mathcal{O}(1)$ where $T_{z}^{\ell^{\prime}}$ is the time $z$ needs to merge itself in level $\ell^{\prime}$. Furthermore, $z$ may be dependent to one of its parents, say $w$. Thus the time to insert $z$ at level $\ell^{\prime}$ is $T_{z}^{\ell^{\prime}}=T_{w}^{\ell^{\prime}}+\mathcal{O}(1)$. If we repeat this argument for all the levels $\hat{\ell}>\ell^{\prime}$ we obtain that the time to merge the cohesive group $X$ at level $\ell^{\prime}$ can be defined as $T_{y}^{\ell^{\prime}}+\ell_{\max}(\mathcal{B})-\ell+1=T_{y}^{\ell^{\prime}}+\ell_{\max}(\mathcal{B})-\ell_{\max}^{\mathcal{B}}(X)+1$ where $y$ is the topmost parent in the pipeline of dependencies. Thus $X$’s overall merge time is $T_{X}^{0}=T_{y}^{0}+\ell_{\max}(\mathcal{B})-\ell_{\max}^{\mathcal{B}}(X)+1=\mathcal{O}(\log n+\ell_{\max}(\mathcal{B})-\ell_{\max}^{\mathcal{B}}(X)+1)$ w.h.p. In the last case, $X$ may depend on its parents for some time and then gain independency. Assume that $X$ depends its parent $z$ until level $\ell^{\prime}$ and than gains independency. $X$’s overall merge time is $T_{X}^{0}=\mathcal{O}(T_{z}^{\ell^{\prime}}+1+\log n)=\mathcal{O}(\log n)$ w.h.p.
@@
Figure 10 shows an example of the WAVE protocol described above, each colored wave represent a round of the merging protocol.
Figure 10: Example of the WAVE protocol. Figure 10(a) and Figure 10(b) show the clean network $\mathcal{C}$ (base skip list) and the preprocessed Buffer network $\mathcal{B}$, respectively. Figure 10(c) shows WAVE’s execution: light gray, light brown, and light blue waves are the first, second and third rounds of the protocol’s execution. After the last round, one step of rewiring is executed to remove the left and right sentries (ls, rs) from $\mathcal{C}$. In this example, the merging phase requires $\mathcal{O}(\log|\mathcal{B}|)$ rounds to merge the buffer network $\mathcal{B}$ with $\mathcal{C}$.

### v2

@@ 2.5 Merge
With all cohesive groups identified, their leaders elected, and dependencies established through parent relationships, the system is now fully prepared to initiate the WAVE protocol for merging into the clean skip list.
The WAVE Protocol
Once the preprocessing completes, the merge proceeds via the WAVE protocol, a top-down, parallel traversal of the clean skip list $\mathcal{C}$ performed by cohesive groups in the buffer $\mathcal{B}$. The protocol resembles a wave-like propagation of merges: each level of $\mathcal{B}$ can become active only after their parents successfully started their merge phase. The WAVE protocol consist of three coordinated behaviors: active traversal, parallel insertion and virtual walking.
Active Traversal.
At the beginning, only the unique top-level cohesive group in $\mathcal{B}$ is active. It initiates the merge by starting at the top-left sentinel of $\mathcal{C}$. The group leader coordinates a rightward traversal in $\mathcal{C}$ to spot the correct insertion point between two existing nodes in $\mathcal{C}$.
As the group moves, each step of the traversal is communicated downwards to its children in $\mathcal{B}$. Specifically, each node in the group broadcast its current position (e.g., the node it has just scanned in $\mathcal{C}$) to its child groups at the levels below. This communication is done in $\mathcal{O}(1)$ rounds because of the preprocessing phase and it provides the children with a “virtual trace” of the parent’s path, allowing them to prepare for future activation as they were walking on $\mathcal{C}$ along with their parents.
If the group encounters a node in $\mathcal{C}$ that triggers a split (e.g., due to key ordering), it divides into subgroups. Each new subgroup elects a new leader and continues the traversal independently.
Parallel Insertion.
Once a cohesive group reaches its insertion point at a given level in $\mathcal{C}$, the leader coordinates a local rewiring operation to insert the group between its predecessors and successors at that level in $\mathcal{C}$. This operation is executed in parallel by all the group members and takes $\mathcal{O}(1)$ rounds. After a successful merge at level $\ell$, the group leader notifies all its children in $\mathcal{B}$. These children are then potentially eligible to activate and begin their own traversal at the next level down.
Virtual Walking.
Nodes in $\mathcal{B}$ that are not yet active remain in an idle state. Each node continuously monitors its two parents at the levels above. As their parents perform their active traversal in $\mathcal{C}$, they broadcast their current search position downward, creating a trace of the search path.
Once both parents complete their merge and communicate readiness, the idle node can change state to active and begins its own merge process.
In addition, by observing these traces, an idle node can determine whether its own search path will likely diverge from those of its parents. If so, the node can safely activate early, using the most recent trace as a starting point. This ensures that the node’s merge proceeds independently, without interfering with its parents’ ongoing insertion.
This mechanism ensures that each level of $\mathcal{B}$ merges as soon as possible only after all dependencies have been satisfied, maintaining a clean, pipelined progression through the levels of the skip list.
The above description outlines the behavior of the WAVE protocol at a high level, highlighting its coordination mechanisms and activation rules. We now present a formal specification of the protocol, broken into two components: (1) Algorithm 5 that describes the behavior of active group leaders as they traverse and insert themselves into the clean skip list $\mathcal{C}$, and (2) Algorithm 6 that highlights the behavior of idle nodes during the virtual walking phase, where they observe parent traces and determine their activation point. Figure 11 shows the Clean Network Buffer $\mathcal{C}$ before the merge phase (Figure 11(a)) and the Buffer Network $\mathcal{B}$ after the preprocessing phase (Figure 11(b)). While Figure 12 provides a conceptual illustration of the WAVE protocol that merges $\mathcal{B}$ into $\mathcal{C}$. At each level, the advancing wave merges the corresponding nodes in $\mathcal{B}$ into $\mathcal{C}$, progressively building the integrated structure top-down. Furthermore, we refer to Appendix D for the visual representation of the step-by-step execution of the protocol on the networks in Figure 11.
Figure 11: Figure 11(a) is the Clean Network $\mathcal{C}$ before the merge phase, while Figure 11(b) shows the Buffer Network $\mathcal{B}$ after the $\mathcal{O}(\log n)$ rounds preprocessing phase.
Figure 12: Conceptual representation of the WAVE protocol. Each wave corresponds to a phase in which cohesive groups at a given level of $\mathcal{B}$ are merged into the clean skip list $\mathcal{C}$, proceeding top-down.
Detailed description of the WAVE Protocol.
In the first round of the merge protocol, the unique cohesive group $X_{\ell_{\max}(\mathcal{B})}=\{x_{0},x_{1},\dots,x_{k}\}$ with $x_{0}=$-$\infty$ and $x_{k}=+\infty$ in $\mathcal{B}$’s topmost level changes its state to merge and starts its merging phase in $\mathcal{C}$ from the left-topmost sentry. For the sake of simplicity, we consider $\mathcal{B}$’s left and right sentries as nodes of the buffer network that must be merged with $\mathcal{C}$ [^9] This does not affect the merge outcome. Moreover, when $\mathcal{B}$ is completely merged with $\mathcal{C}$, we can remove $\mathcal{B}$’s sentinels from $\mathcal{C}$ in $\mathcal{O}(1)$ rounds.. We assume that such sentinels are contained in $\mathcal{C}$’s ones. Moreover, at each round, every cohesive group $X\subseteq\mathcal{B}$ on a node $v\in\mathcal{C}$ at level $\ell$ that is in the merge performs a merge phase step that is composed by a Traversal Step in which the cohesive group moves on $\mathcal{C}$ and (if needed) a Merge Step in which $X$ merges in the current level $\ell$ in the clean network.
Input: Cohesive group $X$ in merge state, leader $x_{0}$
1 while $X$ has not merged at level 0 do
2   Let $\ell\leftarrow$ current level of $X$;
 TraversalStep($x_{0}$, $X$, $\ell$);
    // $\mathcal{O}(1)$ Rounds.
 MergeStep($X$, $\ell$);
    // $\mathcal{O}(1)$ Rounds.
3   Update $X$’s current level to $\ell-1$;
4
5 end while
6Set state of all $u\in X$ to done;
Algorithm 3 Overview of the Merge Phase for Group $X$
Algorithm 3 gives an overview of the step executed by a cohesive group in the merge phase. The group, until merging at level 0, performs a traversal step followed by a potential merge step at the current level. Each of these steps requires $\mathcal{O}(1)$ number of rounds. Algorithm 4 and Algorithm 5 provide a formal description of the steps performed by the nodes while merging. A key idea of the protocol is that every time a node in the merge state performs one move on $\mathcal{C}$, it communicates its progress to all its children. Notice, that because of the preprocessing phase parents can communicate with their children in constant number of rounds. Nodes, that are idle in $\mathcal{B}$ receiving updates from their parents can use this information to “virtually walk” on the clean network along their parents. Indeed, a node $v$, in the buffer network $\mathcal{B}$ in idle state is continuously listening to its parents. $v$’s behavior changes according to the type of message it receives. More precisely, let us assume that $v$ receives a message of the type $m(u)=\langle L,R,\texttt{move},\texttt{level}\rangle$ from its parent $u$: If $v$ is idle and discovers that both its parents successfully merged in $\mathcal{C}$ in level $\ell=\ell_{\max}(v)$, then it must change state to merge and proceed with its own merge phase in $\mathcal{C}$ starting from $v.\texttt{starting\_point}$ at level $v.\texttt{level\_to\_start}$. Algorithm 6 describes the actions performed by the idle nodes in the buffer network upon receiving messages from their parents.
Before analyzing the merge runtime, we first observe that idle nodes only update their state using information from their parents. This ensures the correctness of their updated search path:
Lemma 2.17.
Let $v$ be an idle node in the buffer skip list $\mathcal{B}$ at level $\ell$. Throughout the execution of the WAVE protocol, every update to $v.\texttt{starting\_point}$ results in a node that lies on the optimal search path for $v$ in the augmented skip list $\mathcal{C}^{\prime}=\mathcal{C}\cup\mathcal{B}_{>\ell}$. Where $\mathcal{C}^{\prime}=\mathcal{C}\cup\mathcal{B}_{>\ell}$ is obtained by merging $\mathcal{C}$ and the elements of $\mathcal{B}$ situated in the level above $\ell$.
Proof 2.18.
Each idle node $v$ receives traversal updates exclusively from its two parents in $\mathcal{B}$, which reside at levels strictly greater than $\ell$. These messages reflect nodes visited during the parents’ optimal search paths while merging into $\mathcal{C}$. Since $v$’s own optimal search path in $\mathcal{C}^{\prime}$ would also pass through the same upper-level nodes before descending, any such updated starting_point lies on a valid prefix of $v$’s optimal path. Thus, the virtual walking behavior maintains correctness.
Input: Cohesive group $X$ at level $\ell$ in $\mathcal{B}$, leader $x_{0}$
Output: Traversal continuation or split + merge at level $\ell$
1
@@ 3   Update $X$’s current level to $\ell-1$;
3 $z\leftarrow$ right neighbor of $v$ at level $\ell$ in $\mathcal{C}$;
4 Broadcast $z$ to all followers in $X$;
5
@@ 4 Broadcast $z$ to all followers in $X$;
7 $X_{\texttt{left}}\leftarrow X\setminus X_{\texttt{right}}$;
8
9if $X_{\texttt{right}}\neq\emptyset$ then
10   Elect new leader $x^{\prime}_{0}\leftarrow\min(X_{\texttt{right}})$;
11 foreach $u\in X_{\texttt{right}}$ do
      Send message $\langle v,z,\text{``right''},\ell\rangle$ to $\Gamma(u)$;
       // $\mathcal{O}(1)$ rounds.
12
13   end foreach
14 TraversalStep($x^{\prime}_{0}$, $X_{\texttt{right}}$, $\ell$);
15
16 end if
17
18if $X_{\texttt{left}}\neq\emptyset$ then
19 foreach $u\in X_{\texttt{left}}$ do
      Send message $\langle v,z,\text{``down''},\ell\rangle$ to $\Gamma(u)$ ;
       // $\mathcal{O}(1)$ rounds.
20
21   end foreach
22 MergeStep($X_{\texttt{left}},\ell$);
23
24 end if
Algorithm 4 TraversalStep for Group Leader $x_{0}$
Input: Group $X$ at level $\ell$ in $\mathcal{B}$
Output: Group merged at level $\ell$ and notified children
1
2Insert $X$ between $v$ and $z$ in $\mathcal{C}$ at level $\ell$;
3 foreach $u\in X$ do
4 $y\leftarrow$ right neighbor of $u$ in newly merged segment;
   Send message $\langle u,y,\text{``down''},\ell\rangle$ to $\Gamma(u)$;
    // $\mathcal{O}(1)$ rounds.
   Notify $\Gamma(u)$ of successful merge;
    // $\mathcal{O}(1)$ rounds.
5
6 end foreach
7
8if $\ell>0$ then
9 if $X$ has neighbors in $\mathcal{B}$ at level $\ell$ then
10      Wait for neighbors to be merged at level $\ell-1$;
11
12   end if
13 TraversalStep(leader($X$), $X$, $\ell-1$);
14
15 else
16   Set state of all $u\in X$ to done;
17
18 end if
Algorithm 5 MergeStep for Group $X$
Input: Messages $m(u)$ from parent(s)
Output: Node state update and possible activation
1
2foreach received $m(u)=\langle L,R,\texttt{move},\texttt{level}\rangle$ do
3   Update $v.\texttt{starting\_point}$ and $v.\texttt{level\_to\_start}$ based on $L$, $R$;
4   Update parent locations depending on move;
5
6 end foreach
7
8if $v$ is independent from both parents then
9 if left neighbor $z$ shares same starting point then
10      Set $v$ as follower;
11
12 else
13      Set $v$ as leader;
14      Form group $X$ with $v$ and right neighbors sharing starting point;
15    TraversalStep($v$, $X$, $v.\texttt{level\_to\_start}$);
16
17   end if
18
19 end if
20
21if both parents merged at level $\ell=\ell_{\max}(v)$ then
22   Set $v$ to merge;
23   Proceed with group formation and merging as above;
24
25 end if
Algorithm 6 VirtualWalk for Idle Node $v$
Analysis of the WAVE Protocol
A key aspect of the protocol is that a cohesive group $X$, after being successfully merged at level $\ell$, may need to wait before proceeding to level $\ell-1$. This delay occurs if one of $X$’s neighbors in $\mathcal{C}$ at level $\ell$ is a node from the buffer $\mathcal{B}$ (i.e., one of its parents) that has not yet completed its merge at level $\ell-1$. The following lemma provides an upper bound on the time required for such a group $X$ to fully merge into $\mathcal{C}$, accounting for possible dependencies on its parents at each level.
@@ 23   Proceed with group formation and merging as above;
Assume that the cohesive group $X$ was merged at level $\ell$ of the clean network and has to proceed with its own merge at level $\ell-1$. Let $v$ and $z$ be $X$’s left and right parents in $\mathcal{B}$. We have three cases to consider: (i) $X$ is independent from its parents; (ii) $X$ depends on at least one of its parents through all the levels $\ell^{\prime}<\ell$; and, (iii) $X$ may depend on one of its parents for some time and become independent at some point. In the first case, $X$ is merging itself in a sub-skip list that contains only elements in the clean network. Thus, $X$’s overall merging time requires $\mathcal{O}(\log n)$ w.h.p. To analyze the second case, we notice that as $X$ depends on its parents, also $X$’s parents may depend on their parents as well and so on. We can model the time a cohesive group $X$ at level $\ell$ has to wait in order to start its own merge procedure at level $\ell-1$ as the time it would have taken had it traveled in a pipelined fashion with its parents. Moreover, such pipelining effect can be extended to include, in a recursive fashion, the parents of $X$’s parents, the parents of the parents of $X$’s parents, and so on. Thus, $X$ can be modeled as the tail of such “parents-of-parents” chain (or pipeline), in which $X$ must wait for all the nodes in the pipeline to be merged at level $\ell-1$ before proceeding with its own merge in such level. Moreover, we observe that, by our assumption, $X$ is dependent on at least one of its parents throughout the merge phase. Without loss of generality, let us assume that $X$ depends on one parent, say $z$. In other words, $X$ is never splitting, and its placing itself right after/before $z$ at each level $\ell^{\prime}<\ell$. The time $X$ takes to merge in one level $\ell^{\prime}$ after $z$ is $\mathcal{O}(1)$ (see Lemma 2.12). Thus, $X$ merge itself in some level $\ell^{\prime}$ by time $T_{z}^{\ell^{\prime}}+\mathcal{O}(1)$ where $T_{z}^{\ell^{\prime}}$ is the time $z$ needs to merge itself in level $\ell^{\prime}$. Furthermore, $z$ may be dependent to one of its parents, say $w$. Thus the time to insert $z$ at level $\ell^{\prime}$ is $T_{z}^{\ell^{\prime}}=T_{w}^{\ell^{\prime}}+\mathcal{O}(1)$. If we repeat this argument for all the levels $\hat{\ell}$$\ell^{\prime}$ we obtain that the time to merge the cohesive group $X$ at level $\ell^{\prime}$ can be defined as $T_{y}^{\ell^{\prime}}+\ell_{\max}(\mathcal{B})-\ell+1=T_{y}^{\ell^{\prime}}+\ell_{\max}(\mathcal{B})-\ell_{\max}^{\mathcal{B}}(X)+1$ where $y$ is the topmost parent in the pipeline of dependencies. Thus $X$’s overall merge time is $T_{X}^{0}=T_{y}^{0}+\ell_{\max}(\mathcal{B})-\ell_{\max}^{\mathcal{B}}(X)+1=\mathcal{O}(\log n+\ell_{\max}(\mathcal{B})-\ell_{\max}^{\mathcal{B}}(X)+1)$ w.h.p. In the last case, $X$ may depend on its parents for some time and then gain independency. Assume that $X$ depends its parent $z$ until level $\ell^{\prime}$ and then gains independency. $X$’s overall merge time is $T_{X}^{0}=\mathcal{O}(T_{z}^{\ell^{\prime}}+1+\log n)=\mathcal{O}(\log n)$ w.h.p.
@@
In summary, the WAVE protocol enables a fully distributed and parallel merge of the buffer skip list $\mathcal{B}$ into the clean skip list $\mathcal{C}$, using a top-down wave of coordinated insertions. Cohesive groups ensure local structure and predictability, while the parent-children dependencies across levels enable global synchronization. Through virtual walking and early activation, the protocol supports a high degree of parallelism and completes in $\mathcal{O}(\log n)$ rounds with high probability.